Use the data in the table to calculate the expected activity times, di, and then compute the critical path for each event usi​​​​​​​
Answer #1

Project (a)

For each activity, calculate the expected time u =(a + 4m + b)/6 For each activity, Variance o* =(b-a)/36. Standard deviation

Critical path is the longest path in the network.

It can be identified by 0 slack time.

Critical path is :  1-2-3-4-6-7

Expected time and planned time of project completion = 31.33

Variance = Sum of variances of critical activities = (0.25+0.11+0.69+1+0.11) = 2.17

Standard deviation = sqrt(2.17) = 1.47

z value = (31.33-31.33) / 1.47 = 0

Probability, P(z) = NORMSDIST(0) = 0.5

Probabilities that the events will occur without delay = 0.5